OpenStudy (anonymous):

-x+2y=3 x^2+y=3 How do I solve with substitution?

OpenStudy (anonymous):

Add the 2 equations.

OpenStudy (anonymous):

you probably didn't mean x^2.

OpenStudy (anonymous):

I ment how do I deal with the x squared

OpenStudy (anonymous):

Or if you did, multiply the second equation by -2, and then add.

OpenStudy (anonymous):

-x+2y=3 x^2+y=3 ---> times -2 ---> -2x²-2y=-6 now -x+2y=3 -2x²-2y=-6 ----------

OpenStudy (anonymous):

Isolate one of the variables from one of the equations then sub it into the other.

OpenStudy (anonymous):

basically.

OpenStudy (anonymous):

So can you show us some work?

OpenStudy (anonymous):

@bannannna

OpenStudy (anonymous):

I got to -2x^2-x+3=0

OpenStudy (anonymous):

Yeah usually when you ask Op to show work, they leave. So at that point I leave as well, because I know they aren't willing to learn.

OpenStudy (anonymous):

I didn't leave just app doesn't show me typing

OpenStudy (anonymous):

We have two equations -x + 2y = 3 & x^2 + y = 3 , So you can use any substitution, So here we solve second equation x^2 + y = 3 for y , so we will get y=3 - x^2. Now plug this value y = 3-x^2 in first equation & solve for x, That is - x + 2(3-x^2) =3 , => -x + 6 - 2x^2=3, => -2x^2 -x+3=0 now use quadratic formula & solve for x then plug that value of x in any equation you will get y

OpenStudy (anonymous):

Ohh I see

OpenStudy (anonymous):

I know what to do now thanks

OpenStudy (anonymous):

you are welcome :)